# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ Appendix A List of selected journal papers
[art_128] 2022 Mobility Urban traffic Forecasting
[art_129] 2022 Mobility Flight delay Forecasting
[art_130] 2022 Mobility Urban traffic Forecasting
[art_131] 2023 Mobility Urban traffic Forecasting
[art_132] 2022 Mobility Urban traffic Forecasting
[art_133] 2022 Mobility Bicycle demand Forecasting
[art_135] 2022 Mobility Urban traffic Forecasting
[art_406] 2024 Mobility Multivariate time series Forecasting
[art_408] 2024 Mobility Multivariate time series Forecasting
[art_419] 2024 Mobility Urban traffic Forecasting
[art_420] 2024 Mobility Urban traffic Forecasting
[art_421] 2024 Mobility Metro flow traffic Forecasting
[art_422] 2024 Mobility Urban traffic Forecasting
[art_424] 2024 Mobility Urban traffic Forecasting
[art_425] 2024 Mobility Network traffic prediction Forecasting
[art_427] 2024 Mobility Urban traffic Forecasting
[art_430] 2024 Mobility Urban traffic Forecasting
[art_431] 2024 Mobility Urban traffic Forecasting
[art_432] 2024 Mobility Urban traffic Forecasting
[art_433] 2024 Mobility Urban traffic Forecasting
[art_434] 2024 Mobility Airport performances Forecasting
[art_435] 2024 Mobility Urban traffic Forecasting
[art_438] 2024 Mobility Urban traffic Forecasting
[art_439] 2024 Mobility Urban traffic Forecasting
[art_441] 2024 Mobility Urban traffic Forecasting
[art_442] 2024 Mobility Urban traffic Forecasting
[art_443] 2024 Mobility Flight delay prediction Forecasting
[art_444] 2024 Mobility Urban traffic Forecasting
[art_445] 2024 Mobility Urban traffic Forecasting
[art_446] 2024 Mobility Bike sharing demand Forecasting
[art_447] 2024 Mobility Urban traffic Forecasting
[art_448] 2024 Mobility Urban traffic Forecasting
[art_449] 2024 Mobility Human mobility Forecasting
[art_451] 2024 Mobility Urban traffic Forecasting
[art_452] 2024 Mobility Bus station Profiling
[art_453] 2024 Mobility Urban traffic Forecasting
[art_454] 2024 Mobility Bus stop mobility Forecasting
[art_12] 2023 Other topics Modulation classification Classification
[art_13] 2022 Other topics Modulation classification Classification
[art_118] 2023 Other topics Encrypted traffic classification Classification
[art_136] 2022 Other topics Spatio-temporal events Forecasting
[art_137] 2023 Other topics Quality indicators prediction Forecasting
[art_138] 2023 Other topics Cellular traffic prediction Forecasting
[art_139] 2023 Other topics Seismic data Forecasting
[art_140] 2023 Other topics Structural dynamics Forecasting
[art_141] 2023 Other topics Tailing dam monitoring Forecasting
[art_142] 2023 Other topics Educational videos engagement Forecasting
[art_143] 2023 Other topics Subsurface production Forecasting
[art_144] 2023 Other topics Cellular Network Traffic Forecasting
[art_145] 2023 Other topics Data centers maintenance Forecasting
[art_146] 2023 Other topics Hydraulic runoff Forecasting
[art_147] 2022 Other topics App popularity Forecasting
[art_148] 2021 Other topics Cellular Network Traffic Forecasting
[art_149] 2022 Other topics Water demand Forecasting
[art_151] 2023 Other topics User preference Edge prediction
[art_159] 2021 Other topics Human activity recognition Classification
[art_303] 2023 Other topics Modulation classification Classification
[art_455] 2024 Other topics Polypropylene production data Forecasting
[4] 2024 Other topics Botnet detection Classification
[art_457] 2024 Other topics Cellular KPI prediction Forecasting
[art_458] 2024 Other topics Tungsten floatation plant Forecasting
[art_459] 2024 Other topics Modulation recognition Classification
[art_460] 2024 Other topics IoT intrusion detection Classification
[art_463] 2024 Other topics Modulation recognition Classification
[art_465] 2024 Other topics Crime prediction Forecasting
[art_468] 2024 Other topics Mobile traffic prediction Forecasting
[art_9] 2022 Predictive monitoring Anomaly detection Classification
[art_10] 2023 Predictive monitoring Anomaly detection Forecasting
[art_53] 2023 Predictive monitoring Fault diagnosis Classification
[art_54] 2023 Predictive monitoring Fault diagnosis Classification
[art_55] 2021 Predictive monitoring Fault diagnosis Classification
[art_56] 2023 Predictive monitoring Fault diagnosis Classification
[art_57] 2023 Predictive monitoring Fault diagnosis Classification
[art_58] 2023 Predictive monitoring Fault prediction Forecasting
[art_152] 2023 Predictive monitoring Remaining useful life Forecasting
[art_153] 2023 Predictive monitoring Remaining useful life Forecasting
[art_154] 2023 Predictive monitoring Remaining useful life Forecasting
[art_155] 2021 Predictive monitoring Remaining useful life Forecasting
[art_156] 2022 Predictive monitoring Remaining useful life Forecasting
[art_157] 2022 Predictive monitoring Remaining useful life Forecasting
[art_158] 2023 Predictive monitoring Remaining useful life Forecasting
[art_192] 2023 Predictive monitoring Fault diagnosis Classification
[art_461] 2024 Predictive monitoring Motor temperature Forecasting
[art_469] 2024 Predictive monitoring Fault diagnosis Classification
[art_470] 2024 Predictive monitoring Root cause localization Forecasting
[art_471] 2024 Predictive monitoring Mechanical fault diagnosis Classification
[art_474] 2024 Predictive monitoring Mechanical fault diagnosis Classification
[art_475] 2024 Predictive monitoring Outlier detection Classification
[art_477] 2024 Predictive monitoring Anomaly detection Forecasting
[art_478] 2024 Predictive monitoring Remaining useful life Forecasting
[art_479] 2024 Predictive monitoring Anomaly detection Classification
[art_480] 2024 Predictive monitoring Sensor data Forecasting
[art_481] 2024 Predictive monitoring Anomaly detection Classification
[art_482] 2024 Predictive monitoring Process factors Forecasting
[art_483] 2024 Predictive monitoring Anomaly detection Classification
[art_484] 2024 Predictive monitoring Anomaly detection Classification
[art_486] 2024 Predictive monitoring Fault diagnosis Classification
[art_487] 2024 Predictive monitoring Fault diagnosis Classification
[art_488] 2024 Predictive monitoring Remaining useful life Forecasting
[art_489] 2024 Predictive monitoring Anomaly detection Classification
[art_491] 2024 Predictive monitoring Fault diagnosis Classification

### v4

@@ Appendix A List of selected papers
[84] 2021 Mobility Urban traffic Forecasting
[222] 2021 Mobility Urban traffic Forecasting
[219] 2021 Mobility Urban traffic Forecasting
[428] 2021 Mobility Crowd flow Forecasting
[402] 2021 Mobility Passenger demand Forecasting
[341] 2022 Mobility Urban traffic Forecasting
[276] 2022 Mobility Urban traffic Forecasting
[342] 2022 Mobility Urban traffic Forecasting
[203] 2022 Mobility Urban traffic Forecasting
[259] 2022 Mobility Urban traffic Forecasting
[85] 2022 Mobility Urban traffic Forecasting
[213] 2022 Mobility Urban traffic Forecasting
[370] 2022 Mobility Urban traffic Forecasting
[320] 2022 Mobility Urban traffic Forecasting
[274] 2022 Mobility Urban traffic Forecasting
[211] 2022 Mobility Crowd flow Forecasting
[220] 2022 Mobility Urban traffic Forecasting
[489] 2022 Mobility Connected internet of vehicles Forecasting
[251] 2022 Mobility Taxi demand Forecasting
[490] 2022 Mobility Urban traffic Forecasting
[506] 2022 Mobility Metro passengers Forecasting
[181] 2022 Mobility Urban traffic Forecasting
[278] 2022 Mobility Urban traffic Forecasting
[22] 2022 Mobility Urban traffic Forecasting
[25] 2022 Mobility Flight delay Forecasting
[112] 2022 Mobility Urban traffic Forecasting
[27] 2022 Mobility Urban traffic Forecasting
[71] 2022 Mobility Bicycle demand Forecasting
[169] 2022 Mobility Urban traffic Forecasting
[178] 2023 Mobility Urban traffic Forecasting
[81] 2023 Mobility Urban traffic Forecasting
[524] 2023 Mobility Urban traffic Forecasting
[122] 2023 Mobility Urban traffic Forecasting
[80] 2023 Mobility Urban traffic Forecasting
[422] 2023 Mobility Urban traffic Forecasting
[387] 2023 Mobility Urban traffic congestion Forecasting
[123] 2023 Mobility Visitation patterns Forecasting
[214] 2023 Mobility Urban traffic congestion Classification
[281] 2023 Mobility Urban traffic Forecasting
[145] 2023 Mobility Urban traffic Forecasting
[411] 2023 Mobility Urban traffic Forecasting
[62] 2023 Mobility Urban traffic Forecasting
[218] 2023 Mobility Urban traffic Forecasting
[212] 2023 Mobility Urban traffic Forecasting
[263] 2023 Mobility Urban traffic Forecasting
[262] 2023 Mobility Urban traffic Forecasting
[173] 2023 Mobility Urban traffic Forecasting
[217] 2023 Mobility Urban traffic Forecasting
[200] 2023 Mobility Urban traffic Forecasting
[198] 2023 Mobility Urban traffic Forecasting
[245] 2023 Mobility Bike sharing demand Forecasting
[438] 2023 Mobility Urban traffic Forecasting
[120] 2023 Mobility Urban traffic Forecasting
[111] 2023 Mobility Urban traffic Forecasting
[511] 2023 Mobility Urban traffic Forecasting
[69] 2023 Mobility Urban traffic Forecasting
[171] 2024 Mobility Urban traffic Forecasting
[493] 2024 Mobility Urban traffic Forecasting
[144] 2024 Mobility Urban traffic Forecasting
[95] 2024 Mobility Urban traffic Forecasting
[530] 2024 Mobility Urban traffic Forecasting
[63] 2024 Mobility Urban traffic Forecasting
[523] 2024 Mobility Urban traffic Forecasting
[461] 2024 Mobility Urban traffic Forecasting
[227] 2024 Mobility Urban traffic Forecasting
[32] 2024 Mobility Urban traffic Forecasting
[163] 2024 Mobility Urban traffic Forecasting
[159] 2024 Mobility Urban traffic Forecasting
[385] 2024 Mobility Metro flow traffic Forecasting
[466] 2024 Mobility Urban traffic Forecasting
[407] 2024 Mobility Urban traffic Forecasting
[312] 2024 Mobility Network traffic volume Forecasting
[13] 2024 Mobility Traffic speed Forecasting
[199] 2024 Mobility Traffic speed Forecasting
[330] 2024 Mobility Urban traffic Forecasting
[352] 2024 Mobility Urban traffic Forecasting
[242] 2024 Mobility Urban traffic Forecasting
[420] 2024 Mobility Airport performances Forecasting
[510] 2024 Mobility Urban traffic Forecasting
[406] 2024 Mobility Urban traffic Forecasting
[348] 2024 Mobility Urban traffic Forecasting
[378] 2024 Mobility Urban traffic Forecasting
[419] 2024 Mobility Flight delay prediction Forecasting
[208] 2024 Mobility Traffic speed Forecasting
[250] 2024 Mobility Urban traffic Forecasting
[89] 2024 Mobility Bike sharing demand Forecasting
[126] 2024 Mobility Urban traffic Forecasting
[388] 2024 Mobility Urban traffic Forecasting
[82] 2024 Mobility Urban traffic Forecasting
[193] 2024 Mobility Bus stop mobility pattern Forecasting
[207] 2021 Predictive monitoring Fault diagnosis Classification
[291] 2021 Predictive monitoring Remaining useful life Forecasting
[128] 2022 Predictive monitoring Anomaly detections Forecasting
[221] 2022 Predictive monitoring Anomaly detection Classification
[441] 2022 Predictive monitoring Remaining useful life) Forecasting
[195] 2022 Predictive monitoring Remaining useful life Forecasting
[204] 2023 Predictive monitoring Anomaly detection Classification
[139] 2023 Predictive monitoring Anomaly detection Classification
[226] 2023 Predictive monitoring Anomaly detection Forecasting
[374] 2023 Predictive monitoring Fault diagnosis Classification
[464] 2023 Predictive monitoring Fault diagnosis Classification
[296] 2023 Predictive monitoring Fault diagnosis Classification
[497] 2023 Predictive monitoring Fault prediction Forecasting
[400] 2023 Predictive monitoring Remaining useful life Forecasting
[46] 2023 Predictive monitoring Remaining useful life Forecasting
[382] 2023 Predictive monitoring Remaining useful life Forecasting
[408] 2023 Predictive monitoring Remaining useful life Forecasting
[206] 2023 Predictive monitoring Fault diagnosis Classification
[79] 2024 Predictive monitoring Photovoltaic degradation Forecasting
[272] 2024 Predictive monitoring Anomaly detection Classification
[283] 2024 Predictive monitoring Anomaly detection Classification
[239] 2024 Predictive monitoring Anomaly detection Classification
[31] 2024 Predictive monitoring Motor temperature Forecasting
[196] 2024 Predictive monitoring Fault diagnosis and detection Classification
[86] 2024 Predictive monitoring Root cause localization Forecasting
[389] 2024 Predictive monitoring Fault diagnosis Classification
[96] 2024 Predictive monitoring Outlier detection Classification
[515] 2024 Predictive monitoring Anomaly detection Classification
[488] 2024 Predictive monitoring Remaining useful life Forecasting
[249] 2024 Predictive monitoring Anomaly detection Classification
[224] 2024 Predictive monitoring Sensor data Forecasting
[114] 2024 Predictive monitoring Anomaly detection Classification
[363] 2024 Predictive monitoring Process factors Forecasting
[21] 2024 Predictive monitoring Anomaly detection Classification
[512] 2024 Predictive monitoring Anomaly detection Classification
[379] 2024 Predictive monitoring Fault diagnosis Classification
[476] 2024 Predictive monitoring Fault diagnosis Classification
[194] 2024 Predictive monitoring Remaining useful life Forecasting
[136] 2024 Predictive monitoring Anomaly detection Classification
[381] 2024 Predictive monitoring Fault diagnosis Classification
[252] 2021 Other Cellular traffic Forecasting
[92] 2021 Other Product demand Forecasting
[303] 2021 Other Cellular traffic Forecasting
[459] 2021 Other Cellular Network Traffic Forecasting
[121] 2022 Other Talent demand and supply Forecasting
[435] 2022 Other Modulation classification Classification
[175] 2022 Other Spatial temporal event prediction Forecasting
[501] 2022 Other App popularity Forecasting
[465] 2022 Other Water demand Forecasting
[475] 2023 Other Encrypted traffic Classification
[108] 2023 Other Cellular traffic Forecasting
[380] 2023 Other Modulation classification Classification
[70] 2023 Other Encrypted traffic classification Classification
[492] 2023 Other Quality indicators prediction Forecasting
[404] 2023 Other Cellular traffic prediction Forecasting
[16] 2023 Other Seismic data Forecasting
[410] 2023 Other Structural dynamics Forecasting
[328] 2023 Other Tailing dam monitoring Forecasting
[266] 2023 Other Subsurface production Forecasting
[34] 2023 Other Cellular Network Traffic Forecasting
[344] 2023 Other Data centers maintenance Forecasting
[440] 2023 Other Hydraulic runoff Forecasting
[6] 2023 Other Modulation classification Classification
[150] 2024 Other Anomaly in microservice system Classification
[255] 2024 Other Emergency supply Forecasting
[130] 2024 Other Polypropylene production Forecasting
[3] 2024 Other Botnet detection Classification
[254] 2024 Other Cellular KPI prediction Forecasting
[399] 2024 Other Tungsten floatation plant Forecasting
[487] 2024 Other Modulation recognition Classification
[394] 2024 Other IoT intrusion detection Classification
[454] 2024 Other Modulation recognition Classification
[519] 2024 Other Crime prediction Forecasting
[170] 2024 Other Mobile traffic prediction Forecasting
[450] 2020 Generic Multivariate time series Forecasting
[277] 2020 Generic Multivariate time series Forecasting
[424] 2020 Generic Multivariate time series Forecasting
[26] 2020 Generic Multivariate time series Forecasting
[338] 2021 Generic Multivariate time series Forecasting
[61] 2021 Generic Multivariate time series Forecasting
[50] 2021 Generic Multivariate time series Forecasting
[356] 2021 Generic Multivariate time series Forecasting
[327] 2021 Generic Multivariate time series Forecasting
[289] 2021 Generic Multivariate time series Classification
[155] 2021 Generic Generic Both
[499] 2022 Generic Multivariate time series Forecasting
[448] 2022 Generic Multivariate time series Forecasting
[51] 2022 Generic Time series Forecasting
[271] 2022 Generic Multivariate time series Forecasting
[48] 2022 Generic Multivariate time series Forecasting
[268] 2022 Generic Multivariate time series Forecasting
[472] 2022 Generic Multivariate time series Classification
[47] 2022 Generic Multivariate time series Forecasting
[52] 2022 Generic Multivariate time series Forecasting
[397] 2022 Generic Multivariate time series Forecasting
[138] 2022 Generic Multivariate time series Forecasting
[83] 2023 Generic Multivariate time series Both
[153] 2023 Generic Multivariate time series Forecasting
[451] 2023 Generic Multivariate time series Forecasting
[59] 2023 Generic Multivariate time series Forecasting
[409] 2023 Generic Multivariate time series Forecasting
[369] 2023 Generic Multivariate time series Forecasting
[49] 2023 Generic Multivariate time series Forecasting
[368] 2023 Generic Multivariate time series Forecasting
[300] 2023 Generic Multivariate time series Forecasting
[36] 2023 Generic Batch workloads Forecasting
[232] 2023 Generic Node classification for websites Classification
[129] 2023 Generic Clinical risk classification Classification
[60] 2023 Generic Sensor network Forecasting
[241] 2023 Generic Multivariate time series Forecasting
[99] 2023 Generic Multivariate time series Forecasting
[177] 2023 Generic Multivariate time series Forecasting
[504] 2023 Generic Long term series forecasting Forecasting
[76] 2023 Generic Multivariate time series Forecasting
[243] 2023 Generic Multivariate time series Forecasting
[118] 2023 Generic Multivariate time series Forecasting
[40] 2023 Generic Multivariate time series Forecasting
[287] 2023 Generic Multivariate time series Forecasting
[395] 2023 Generic Multivariate time series Forecasting
[355] 2023 Generic Multivariate time series Forecasting
[186] 2023 Generic Multivariate time series Forecasting
[473] 2023 Generic Multivariate time series Forecasting
[57] 2023 Generic Multivariate time series Forecasting
[15] 2023 Generic Impact of graph construction Forecasting
[432] 2024 Generic Multivariate time series Forecasting
[458] 2024 Generic Multivariate time series Forecasting
[436] 2024 Generic Multivariate time series Forecasting
[286] 2024 Generic Multivariate time series Forecasting
[491] 2024 Generic Multivariate time series Forecasting
[279] 2024 Generic Multivariate time series Forecasting
[456] 2024 Generic Multivariate time series Classification
[143] 2024 Generic Multivariate time series Forecasting
[505] 2024 Generic Multivariate time series Forecasting
[244] 2024 Generic Multivariate time series Forecasting
[74] 2024 Generic Multivariate time series Classification
[113] 2024 Generic Multivariate time series Classification
[396] 2024 Generic Multivariate time series Both
[125] 2024 Generic Multivariate time series Forecasting
[261] 2024 Generic Multivariate time series Classification
[231] 2024 Generic Multivariate time series Forecasting
[100] 2024 Generic Multivariate time series Classification
[137] 2024 Generic Multivariate time series Forecasting
[152] 2024 Generic Multivariate time series Forecasting
[453] 2024 Generic Multivariate time series Forecasting
[390] 2024 Generic Multivariate time series Forecasting
[360] 2024 Generic Multivariate time series Forecasting
[401] 2024 Generic Multivariate time series Both
[337] 2024 Generic Multivariate time series Forecasting
[240] 2024 Generic Multivariate time series Forecasting
[140] 2024 Generic Multivariate time series Classification
[443] 2024 Generic Multivariate time series Forecasting